if a delivery service delivered to all points within 25 miles. if they went 20 miles east and 14 miles south will there be a delivery charge. can i get the solution in simple terms for this question.

Answer by rwm(914)

This needs the Pythagorean theorem.20^2 + 14^2=c^2
400+196=596
sqrt(596)= 24.4